Check my work Be sure to answer all parts. Earths oceans have an average depth of 3800 m, a total area of 3.63 x 103 km, and an average concentration of dissolved gold of 5.8 x 10L. If a recent price of gold was $757.00/troy oz, what is the value of gold in the oceans? (I troy oz-31.1 g) Enter your answer in scientific notation.
Answer #1

--Average depth of an ocean = 3800 m = 3.8 x 103 m

--Total surface area of ocean surface = 3.63 X 103 km2 [1km = 102 m,1km2 = (103m)2 = 106 m2]

= 3.63 X 103 X 106 m2

= 3.63 X 109m2

--Therefore, Volume of the ocean water = Surface area of ocean X depth of the ocean

= (3.63 X 109m2 ) x (3.8 x 103 m)

= 13.794 X 1012m3    [1L=0.001 m3 = 10-3m3]

13.794 x 1012m3 10-3,n3LT

= 13.794 X 1015 L

= 1.3794 x 1016 L

-- Concentration of dissolved gold in ocean water = 5.8 X 10-9 gL-1

Amount of gold dissolved in 1.3794 x 1016 L of ocean water = (5.8 X 10-9 gL-1) X 1.3794 x 1016 L

= 8.00052 x 107 g

-- Amount of gold interms of troy oz =

  = \frac{8.00052\times 10^{7}g}{31.1g/troy oz} = 0.2573 \times 10^{7} troy oz

= 2.573 X 106 troy oz

--Recent prize of the gold is = $ 757.0/troy oz

The value of the gold present in ocean = (2.573 X 106 troy oz) X ($ 757.0/troy oz)

= $ 1947.3934 X 106

= $ 1.9474 X 109

The value of the gold present in the ocean of depth 3.8 x 103 m and surface area 3.63 X 103 km2 is

= $ 1.9474 X 109